35 heads in the cage and 94 feet.
How many rabbits & pheasants?...
How do you work out this equation please, thank you.

Answer by stanbon(75887)

35 heads in the cage and 94 feet.
How many rabbits & pheasants?...
---------------------
Head Equation: r + p = 35
Feet Equation: 4r+2p = 94
-----------------------------
Modify:
2r + 2p = 70
4r + 2p = 94
--------
Solve for "r":
2r = 24
r = 12 (# of rabbits)
------------
Solve for "p":
r + p = 35
12 + p = 35
p = 23 (# of pheasants)
===============
Cheers,
Stan H.
===============